Write a system of equations that has the point (3,2) as the only solution. How do you do that?
Answer by user_dude2008(1825)

(3,2) ---> x=3, y=2

Note: 3+2=5 and 3-2=1

One system:

x+y=5
x-y=1


Check by plugging in x=3, y=2